Calculate the mean free path of a gas of randomly-moving hard spheres of 2 nanometer radius, when the density of spheres is 3.00 x 1019 per cubic meter.


  • A. 1.17 × 10-4 m
  •  B. 9.76 × 10-5 m
  •  C. 4.70 × 10-4 m
  •  D. 1.88 × 10-3 m
  •  E. None of the above.
1
Expert's answer
2021-06-04T08:05:40-0400

The mean free path or the average distance traveled between collisions can be calculated by using


"\lambda =v\cdot t_{mean}= \frac{V}{4\pi \sqrt{2}r²N}= \frac{1}{4\pi \sqrt{2}r² \rho}"


Then, we substitute the density of molecules "(\rho=N/V)" and the molecule radius r (in meters) to find "\lambda" :


"\lambda = \frac{1}{4\pi \sqrt{2}r² \rho}= \frac{1}{4\pi \sqrt{2}(2\times 10^{-9}m)² (3\times10^{19}\,molecules/m^{3})}=4.67\times10^{-4}\,m/molecule"


In conclusion, the correct answer is C. 4.70 × 10-4 m


Reference:

  • Young, H. D., Freedman, R. A., & Ford, A. L. (2006). Sears and Zemansky's university physics (Vol. 1). Pearson education.